### v1

# Flows of linear orders on sparse graphs

Abstract.
We consider the topological dynamics of the automorphism group of a particular sparse graph $M_{1}$ resulting from an ab initio Hrushovski construction. We show that minimal subflows of the flow of linear orders on $M_{1}$ have all orbits meagre, partially answering a question of Tsankov regarding results of Evans, Hubička and Nešetřil on the topological dynamics of automorphism groups of sparse graphs.
Key words and phrases:
sparse graphs, Hrushovski constructions, admissible orders, meagre orbits, orientations
2020 Mathematics Subject Classification:
03C15, 37B05, 20B27, 05C55, 05D10
This project formed the second part of the PhD of the author at Imperial College London, under the supervision of Prof David Evans.
1. Introduction
The paper [13] of Kechris, Pestov and Todorčević established links between topological dynamics and structural Ramsey theory, with further developments in [17], [19], [2] (among others). We assume the reader is familiar with the background here, and briefly recall three key results, which we formulate for strong classes (classes of structures where we restrict to a particular subclass of permitted embeddings – see [4, Definition 2.1]):
Theorem ([2, Theorems 1.1 & 1.2, Corollary 3.3]).
Let $G$ be a Polish group with universal minimal flow $M(G)$.
  1. (1)
    $M(G)$ is metrisable iff $G$ has a coprecompact extremely amenable closed subgroup;
  2. (2)
    if $M(G)$ is metrisable, then $M(G)$ has a comeagre orbit.
Theorem ([13, Theorem 4.8]).
Let $M$ be a Fraïssé limit of a strong amalgamation class $(\mathcal{K},\leq)$. Then $\operatorname{Aut}(M)$ is extremely amenable iff $(\mathcal{K},\leq)$ is a Ramsey class of rigid structures.
Theorem ([13, Theorem 10.8], [17, Theorem 5], [19, Theorem 5.7]).
Let $M$ be the Fraïssé limit of an amalgamation class $(\mathcal{K},\leq)$, and let $N$ be the Fraïssé limit of an amalgamation class $(\mathcal{K}^{+},\leq^{+})$ of rigid structures which is a reasonable strong expansion of $(\mathcal{K},\leq)$. Let $G=\operatorname{Aut}(M),H=\operatorname{Aut}(N)$. Suppose $(\mathcal{K}^{+},\leq^{+})$ has the Ramsey property and the expansion property over $(\mathcal{K},\leq)$, and suppose $H$ is a coprecompact subgroup of $G$.
Then the universal minimal flow $M(G)$ of $G$ is metrisable and has a comeagre orbit. Explicitly, we have $M(G)=\widehat{G/H}$, the completion of the quotient $G/H$ of the right uniformity on $G$.
(We can also describe the comeagre orbit explicitly – see the references for further details.)
The paper [4], which was the starting point for the current paper, showed that classes of sparse graphs used in Hrushovski constructions ([10], [11]) demonstrate different behaviour to classes previously studied in the KPT context. A graph $A$ is $k$-sparse if for all finite $B\subseteq A$, we have $|E(B)|\leq k|B|$. It is well-knwon ([15]) that a graph is $k$-sparse iff it is $k$-orientable. We take $k=2$ for presentational simplicity.
We briefly describe the classes $\mathcal{C}_{0},\mathcal{C}_{1},\mathcal{C}_{F}$ of sparse graphs found in [4]. Let $\mathcal{C}_{0}$ denote the class of finite $2$-sparse graphs. For $A,B\in\mathcal{C}_{0}$, we write $A\leq_{s}B$ if there exists a $2$-orientation of $B$ in which $A$ is successor-closed (by [6, Lemma 1.5], this is equivalent to another phrasing in terms of predimension). With this notion of $\leq_{s}$-substructure, we have the free amalgamation class $(\mathcal{C}_{0},\leq_{s})$ with Fraïssé limit $M_{0}$ (this structure, an “ab initio Hrushovski construction”, was first described in [11]).
We also have a “simplified” version of $M_{0}$, first considered in [6]. Let $\mathcal{D}_{1}$ denote the class of finite $2$-oriented graphs with no directed cycles, and let $\mathcal{C}_{1}$ be the class of graph reducts of structures in $\mathcal{D}_{1}$. For $A,B\in\mathcal{C}_{1}$, write $A\leq_{1}B$ if there exists an expansion $B^{+}\in\mathcal{D}_{1}$ in which $A$ is successor-closed. Then $(\mathcal{C}_{1},\leq_{1})$ is again a free amalgamation class, whose Fraïssé limit we denote by $M_{1}$.
The structures $M_{0},M_{1}$ are not $\omega$-categorical, but if we consider $2$-sparse graphs whose predimension is greater than a certain control function $F$ with logarithmic growth, and take another notion of $\leq_{d}$-substructure (where the predimension strictly increases), we obtain an $\omega$-categorical Fraïssé limit $M_{F}$ (see [4] and [3, Section 3] for details).
We then have:
Theorem ([4, Theorems 3.7, 3.16]).
Let $M=M_{1},M_{0},M_{F}$. Then $\operatorname{Aut}(M)$ has no coprecompact extremely amenable closed subgroup, and so its universal minimal flow is non-metrisable.
Equivalently, by [13, Theorem 4.8], we have that $M$ has no coprecompact Ramsey expansion.
The case of $M_{F}$ is particularly interesting as it shows that the automorphism group of an $\omega$-categorical structure need not have “tame” dynamics in the sense of metrisability of the universal minimal flow, and that $\omega$-categorical structures are not necessarily tame from a structural Ramsey theory perspective either.
The paper [4] also investigates the existence of comeagre orbits. Let $\operatorname{Or}(M)$ denote the $\operatorname{Aut}(M)$-flow of $2$-orientations on $M$.
Theorem ([4, Theorem 5.2]).
Let $M=M_{1},M_{0},M_{F}$. Let $Y$ be a minimal subflow of $\operatorname{Or}(M)$. Then all $\operatorname{Aut}(M)$-orbits of $Y$ are meagre.
Note that if $M(G)$ has a comeagre orbit, then so does any minimal $G$-flow (see [1]), so the above result shows again that $\operatorname{Aut}(M)$ for $M=M_{1},M_{0},M_{F}$ has non-metrisable universal minimal flow, using [2, Theorem 1.2]. In the context of the above result, T. Tsankov asked the following ([4, concluding remarks]):
Question (Tsankov).
Let $M=M_{1},M_{0},M_{F}$. Does $\operatorname{Aut}(M)$ have a (non-trivial) metrisable minimal flow with a comeagre orbit?
David Evans suggested that the author investigate the $\operatorname{Aut}(M)$-flow $\mathcal{LO}(M)$ of linear orders on $M$. We obtain the following result (the main result of this paper), for $M_{1}$, the “simplified version” of $M_{0}$ (see 2.4):
Theorem 4.1.
Let $Y\subseteq\mathcal{LO}(M_{1})$ be a minimal subflow. Then all $\operatorname{Aut}(M_{1})$-orbits on $Y$ are meagre.
This result demonstrates that the phenomenon seen in [4, Theorem 5.2] occurs more generally for other flows on $M_{1}$, partially answering the question of Tsankov. We also note in passing that $M_{1}$ is $\omega$-saturated and its theory is $\omega$-stable (see [6]).
To prove 4.1, we take the class of finite ordered graphs which $\leq$-embed into some element of the minimal flow $Y$, and show that this class fails to have the weak amalgamation property – this gives the result, using 2.32. To show failure of the weak amalgamation property, we will use the Ramsey expansion given by the admissible orders, from [5, Section 3.1]. We discuss these in Section 3.
The author has not been able to extend Theorem 4.1 to $M_{0}$ and $M_{F}$, and believes that the proof strategy for $M_{1}$ would require significant modification for these cases. Partial results for $M_{0}$ (giving some information about minimal subflows of $\mathcal{LO}(M_{0})$, and clarifying obstructions to the proof strategy) can be found in Chapter 5 of [18], the author’s PhD thesis.
Recall that we consider $M_{1}$ to be a “simplified version” of $M_{0}$. It would be interesting to know if there is an analogous “simplified version” of $M_{F}$ – if so, it may be possible to prove 4.1 for an $\omega$-categorical structure. See [18, Chapter 7].
2. Background
We briefly summarise the background required for Section 3 and 4.1. This section is mostly based on [4, Section 2].
All first-order languages considered in this paper will be countable and relational, unless specified otherwise. (For the Ramsey result that we use, we will also need to consider countable languages consisting of relation symbols and set-valued function symbols, as in [5].)
2.1. Graphs and oriented graphs: notation and setup
A graph $(A,\operatorname{E}(A))$ consists of a vertex set $A$ and an edge set $\operatorname{E}(A)\subseteq A^{(2)}$ (where $A^{(2)}$ denotes the set of unordered pairs). Here we only work with simple graphs: graphs with no loops or multiple edges. A set $\rho_{A}\subseteq A^{2}$ is an orientation of $(A,\operatorname{E}(A))$ if:
  • for $xy\in\operatorname{E}(A)$, exactly one of $(x,y),(y,x)$ is in $\rho_{A}$;
  • for $(x,y)\in\rho_{A}$, $xy\in\operatorname{E}(A)$.
We refer to $(A,\operatorname{E}(A),\rho_{A})$ as an oriented graph. We usually just write $(A,\rho_{A})$ or $A$, the edge set being clear from context.
We work with graphs in first-order languages as usual: the language of graphs $L_{\text{graph}}=\{E\}$ with $E$ binary, and a $L_{\text{graph}}$-structure $(A,E_{A})$ is then a graph if $E_{A}\subseteq A^{2}$ is symmetric and irreflexive. We expand $L_{\text{graph}}$ to the language of oriented graphs $L_{\text{or}}=\{E,\rho\}$, where $\rho$ is binary. A $L_{\text{or}}$-structure $(A,E_{A},\rho_{A})$ is an oriented graph if $\rho_{A}$ is an orientation of the graph $(A,E(A))$, where $E(A)$ is the edge set of unordered pairs induced by $E_{A}$. When we refer to a subgraph of a graph, or an oriented subgraph of an oriented graph, we mean an $L_{\text{graph}}$- or $L_{\text{or}}$-substructure respectively.
Definition 2.1.
Let $(A,\rho_{A})$ be an oriented graph. If $(x,y)\in\rho_{A}$, we write $xy\in\rho_{A}$. We call $xy$ an out-edge of $x$ and an in-edge of $y$. We call $y$ an out-vertex or successor of $x$, and $x$ an in-vertex or predecessor of $y$. The out-neighbourhood $\operatorname{N_{{}_{\!+}}}(x)$ and in-neighbourhood $\operatorname{N_{{}_{\!--}}}(x)$ of $x$ consist of the out-vertices and in-vertices respectively of $x$, and we define the out-degree $\operatorname{d_{{}_{+}}\!}(x)=|\operatorname{N_{{}_{\!+}}}(x)|$ and the in-degree $\operatorname{d_{{}_{--}}\!}(x)=|\operatorname{N_{{}_{\!--}}}(x)|$.
2.2. Sparse graphs: $\mathcal{C}_{0}$ and $\mathcal{C}_{1}$
Definition 2.2.
Let $k\in\mathbb{N}_{+}$. A graph $A$ is $k$-sparse if for all $B\subseteq_{\text{fin.\!}}A$, $|\operatorname{E}(B)|\leq k|B|$.
Let $(A,\rho_{A})$ be an oriented graph. We call $\rho_{A}$ a $k$-orientation if for $x\in A$, $|\operatorname{N_{{}_{\!+}}}(x)|\leq k$. If an undirected graph $A$ has a $k$-orientation, we say it is $k$-orientable.
The following well-known proposition ([15]) will be a key tool here.
Proposition 2.3 ([4], Th. 3.4).
A graph $A$ is $k$-orientable iff it is $k$-sparse.
For presentational simplicity, we will usually work with $k=2$. Our results generalise straightforwardly for $k>2$.
Definition 2.4.
Let $\mathcal{C}_{0}$ be the class of finite $2$-sparse graphs, and let $\mathcal{D}_{0}$ be the class of finite $2$-oriented graphs. By Proposition 2.3 we have that $\mathcal{C}_{0}$ is the class of graph reducts of $\mathcal{D}_{0}$.
Let $\mathcal{D}_{1}$ be the class of finite $2$-oriented graphs with no directed cycles. By a slight abuse of terminology, we will refer to a $2$-oriented graph with no directed cycles as an acyclic $2$-oriented graph. Let $\mathcal{C}_{1}$ be the class of graph reducts of $\mathcal{D}_{1}$.
We may also define $\mathcal{C}_{1}$ directly, as per the lemma below.
Lemma 2.5 ([6], Lem. 1.3).
A finite graph $A$ has a $k$-orientation with no directed cycles iff every non-empty subgraph $B$ has a vertex of degree $\leq k$ in $B$.
Proof.
:absent:$\Rightarrow$:⇒ : Consider the induced orientation on $B$ from an acyclic $k$-orientation on $A$. As this orientation is acyclic, $B$ contains a vertex $v$ which has no in-edges in $B$, and therefore the degree of $v$ in $B$ is equal to its out-degree in $B$, which is $\leq k$.
:absent:$\Leftarrow$:⇐ : We use induction on $|A|$. For $|A|=1$, the claim is trivial. For the inductive step, take a vertex $a\in A$ of degree $\leq k$. By assumption, we may give $A\setminus\{a\}$ an acyclic $k$-orientation. Then orient the edges of $a$ outwards from $a$ to produce an acyclic $k$-orientation of $A$. ∎
Thus $\mathcal{C}_{1}$ consists of the finite graphs $A$ where every non-empty subgraph $B\subseteq A$ has a vertex of degree $\leq 2$.
2.3. Strong classes
To produce amalgamation classes from the classes of structures defined in the previous section, we will consider certain distinguished subclasses of embeddings. We assume the reader is familiar with Fraïssé theory for strong classes: see [3, Section 3] for more details.
Definition 2.6.
Let $\mathcal{K}$ be a class of finite $L$-structures closed under isomorphisms. Let $\mathcal{S}\subseteq\operatorname{Emb}(\mathcal{K})$ be a class of embeddings between structures in $\mathcal{K}$ which contains all isomorphisms and is closed under composition, and which also satisfies:
  1. ()($\ast$)( ∗ )
    if $f:A\to C$ is in $\mathcal{S}$ and $f(A)\subseteq B\subseteq C$ with $B\in\mathcal{K}$, then $f:A\to B$ is in $\mathcal{S}$.
Then we call $(\mathcal{K},\mathcal{S})$ a strong class, and call the elements of $\mathcal{S}$ strong embeddings.
If $A,B\in\mathcal{K}$, $A\subseteq B$ and the inclusion map $\iota:A\hookrightarrow B$ is in $\mathcal{S}$, then we write $A\leq B$ and say $A$ is a strong substructure of $B$. We have that $\leq$ is reflexive and transitive, and if $A\leq C$ and $A\subseteq B\subseteq C$ with $B\in\mathcal{K}$, then $A\leq B$. We will often write $(\mathcal{K},\leq)$ instead of $(\mathcal{K},\mathcal{S})$, and we will use different symbols resembling $\leq$ (e.g. $\leq_{1},\leq_{s},\sqsubseteq_{s}$) to indicate different classes of embeddings.
Let $(\mathcal{K},\leq)$ be a strong class. Then we have that for $f:A\to B$ in $\mathcal{S}$, if $X\leq A$, then $f(X)\leq B$.
We now define strong substructures of infinite structures. Let $(\mathcal{K},\leq)$ be a strong class. Let $A_{1}\leq A_{2}\leq\cdots$ be an increasing $\leq$-chain of structures in $\mathcal{K}$ with union $M$. Let $A\subseteq_{\text{fin.\!}}M$. Then we write $A\leq M$ to mean that there is some $A_{i}$ ($i\geq 1$) with $A\leq A_{i}$. By ()($\ast$)( ∗ ), this definition is independent of the particular $\leq$-chain we consider.
Let $g\in\operatorname{Aut}(M)$. It is easy to see that if $A\leq M$, then $gA\leq M$: that is, all $g\in\operatorname{Aut}(M)$ preserve $\leq$.
2.4. Fraïssé theory for strong classes
The classical Fraïssé theory directly generalises to strong classes. (See [3, Section 3]. For the classical Fraïssé theory, originally developed in [8], see [9].)
Given a strong class $(\mathcal{K},\leq)$, we define the joint embedding property (JEP), amalgamation property (AP) and free amalgamation as before, but we restrict to $\leq$-embeddings.
Definition 2.7.
Let $(\mathcal{K},\leq)$ be a strong class, and let $M$ be a union of an increasing $\leq$-chain of structures in $\mathcal{K}$. Let $\operatorname{Age}_{\leq}(M)$ denote the class of finite structures which $\leq$-embed into $M$.
$M$ has the $\leq$-extension property if for all $A,B\in\operatorname{Age}_{\leq}(M)$ and $\leq$-embeddings $f:A\to M,g:A\to B$, there exists a $\leq$-embedding $h:B\to M$ with $hg=f$.
$M$ is $\leq$-homogeneous if each isomorphism $f:A\to A^{\prime}$ between finite $A,A^{\prime}\leq M$ extends to an automorphism of $M$.
Definition 2.8.
Let $(\mathcal{K},\leq)$ be a strong class. We say that $(\mathcal{K},\leq)$ is an amalgamation class or Fraïssé class if $(\mathcal{K}\,\leq)$ contains countably many isomorphism types, and has the joint embedding and amalgamation properties.
Theorem 2.9 (Fraïssé-Hrushovski).
Let $(\mathcal{K},\leq)$ be an amalgamation class. Then there is a structure $M$ which is a union of an increasing $\leq$-chain, such that $M$ is $\leq$-homogeneous and $\operatorname{Age}_{\leq}(M)=\mathcal{K}$. Such an $M$ is unique up to isomorphism.
We call this structure the Fraïssé limit of $\mathcal{K}$.
2.5. Strong expansions
We will often have the situation where we have a strong class of $L$-structures together with a strong class of $L^{+}$-structures (with a potentially different notion of distinguished embedding), where $L^{+}\supseteq L$ is a relational language expanded from $L$.
Definition 2.10 ([4], Def. 2.9).
Let $(\mathcal{K},\leq)$ be a strong class of $L$-structures, and let $(\mathcal{K}^{+},\leq^{+})$ be a strong class of $L^{+}$-structures. We call $(\mathcal{K}^{+},\leq^{+})$ is a strong expansion of $(\mathcal{K},\leq)$ if:
  1. (1)
    $\mathcal{K}$ is the class of $L$-reducts of $\mathcal{K}^{+}$;
  2. (2)
    for $\leq^{+}$-strong $f:A^{+}\to B^{+}$, $f:A^{+}|_{L}\to B^{+}|_{L}$ is $\leq$-strong;
  3. (3)
    for $\leq$-strong $f:A\to B$ and $A^{+}\in\mathcal{K}^{+}$ an expansion of $A$, there exists an expansion $B^{+}\in\mathcal{K}^{+}$ of $B$ such that $f:A^{+}\to B^{+}$ is $\leq^{+}$-strong.
Lemma 2.11.
Let $(\mathcal{K}^{+},\leq^{+})$ be a strong expansion of the strong class $(\mathcal{K},\leq)$. If $(\mathcal{K}^{+},\leq^{+})$ is a (free) amalgamation class, then so is $(\mathcal{K},\leq)$. If $M^{+}$ is the Fraïssé limit of $(\mathcal{K}^{+},\leq^{+})$, then $M=M^{+}|_{L}$ is the Fraïssé limit of $(\mathcal{K},\leq)$, and $\operatorname{Aut}(M^{+})$ is a closed subgroup of $\operatorname{Aut}(M)$.
The proof is straightforward.
Definition 2.12.
A particular case of Definition 2.10 is the order expansion of a strong class. Let $(\mathcal{K},\leq)$ be a strong class of $L$-structures, and let $L^{\prec}=L\cup\{\prec\}$ with $\prec$ binary. Let $\mathcal{K}^{\prec}$ be the class of $L^{+}$-structures $(A,\prec_{A})$, where $A\in\mathcal{K}$ and $\prec_{A}$ is a linear order on $A$. For $(A,\prec_{A}),(B,\prec_{B})\in\mathcal{K}^{\prec}$, write $(A,\prec_{A})\leq(B,\prec_{B})$ if $A\leq B$ and $\prec_{A}=\prec_{B}|_{A}$. We then have that $(\mathcal{K}^{\prec},\leq)$ is a strong class and a strong expansion of $(\mathcal{K},\leq)$. We call $(\mathcal{K}^{\prec},\leq)$ the order expansion of the strong class $(\mathcal{K},\leq)$.
2.6. Sparse graphs: $\sqsubseteq_{s}$ and $\leq_{1}$
We will now describe the distinguished notions of embedding involved in defining particular strong classes for $\mathcal{C}_{1}$ and $\mathcal{D}_{1}$. (The material in this section is based on Section 3.4 of [4], and for $\leq_{1}$ was originally developed in [7].)
Definition 2.13.
Let $A$ be an oriented graph. For $B\subseteq A$, we write $B\sqsubseteq_{s}A$ to mean that $B$ is successor-closed in $A$: for $b\in B$, if $ba\in\rho_{A}$ then $a\in B$. Note that intersections of successor-closed subsets are also successor-closed. For $B\subseteq A$, the successor-closure $\operatorname{scl}(B)$ is the smallest successor-closed subset of $A$ containing $B$.
Definition 2.14.
Let $A,B\in\mathcal{C}_{1}$ with $A\subseteq B$. We write $A\leq_{1}B$ if there exists an acyclic $2$-orientation $B^{+}\in\mathcal{D}_{1}$ of $B$ in which the induced orientation $A^{+}\in\mathcal{D}_{1}$ on $A$ has $A^{+}\sqsubseteq_{s}B^{+}$.
Lemma 2.15.
Let $B\in\mathcal{C}_{1}$.
  1. (1)
    $A\leq_{1}B$, $X\subseteq B$Rightarrow $A\cap X\leq_{1}X$.
  2. (2)
    $A\leq_{1}C\leq_{1}B$Rightarrow $A\leq_{1}B$.
  3. (3)
    $A_{1},A_{2}\leq_{1}B$Rightarrow $A_{1}\cap A_{2}\leq_{1}B$.
The above three statements also hold for $B\in\mathcal{D}_{1}$ and $\sqsubseteq_{s}$.
Proof.
(1): immediate. (2): let $B^{+}$ be an acyclic $2$-orientation of $B$ in which $C$ is successor-closed. Create a new acyclic $2$-orientation $B^{++}$ on $B$ by taking $B^{+}$ and replacing the induced orientation on $C$ with an acyclic $2$-orientation on $C$ in which $A$ is successor-closed. Then $A$ is successor-closed in $B^{++}$. (3): follows from (1) and (2).
For $B\in\mathcal{D}_{1}$ and $\sqsubseteq_{s}$, the proof of the above three statements is straightforward. ∎
By the above lemma (and similar arguments to the proof of (2) in Lemma 2.15), we have that:
Lemma 2.16.
$(\mathcal{C}_{1},\leq_{1})$ and $(\mathcal{D}_{1},\sqsubseteq_{s})$ are strong classes with free amalgamation, and $(\mathcal{D}_{1},\sqsubseteq_{s})$ is a strong expansion of $(\mathcal{C}_{1},\leq_{1})$.
Let $M_{1}$ be the Fraïssé limit of $(\mathcal{C}_{1},\leq_{1})$ and let $G_{1}=\operatorname{Aut}(M_{1})$.
2.7. Topological dynamics and Ramsey classes
A $G$-flow is a continuous action $G\curvearrowright X$ of a Hausdorff topological group $G$ on a nonempty compact Hausdorff topological space $X$. We will often simply write $X$ to refer to the $G$-flow $G\curvearrowright X$ when this is clear from context. Given $G\curvearrowright X$, the orbit closure $\overline{G\cdot x}$ of a point $x\in X$ is a $G$-invariant compact subset of $X$. In general, a nonempty compact $G$-invariant subset $Y\subseteq X$ defines a subflow by restricting the $G$-action to $Y$. A $G$-flow on $X$ is minimal if it contains no proper subflows. A $G$-flow is minimal iff every orbit is dense. By Zorn’s lemma, every $G$-flow contains a minimal subflow.
Let $X,Y$ be $G$-flows. A $G$-flow morphism $X\to Y$ is a continuous map $\pi:X\to Y$ such that $\pi(g\cdot x)=g\cdot\pi(x)$ (this property is called $G$-equivariance). Bijective $G$-flow morphisms are isomorphisms, as they are between compact Hausdorff spaces. If $Y$ is minimal, then any $G$-flow morphism $X\to Y$ is surjective, as the image is a subflow.
Definition 2.17.
Let $G$ be a Hausdorff topological group. If every $G$-flow has a $G$-fixed point, we call $G$ extremely amenable.
We now define the Ramsey property for strong classes. (This section is based on Section 2.2 of [4].)
Definition 2.18.
Let $(\mathcal{K},\leq)$ be a strong class. For $A,B\in\mathcal{K}$, we write $\binom{B}{A}=\{A^{\prime}\leq B:A^{\prime}\cong A\}$ for the set of $\leq$-copies of $A$ inside $B$. (Note that by the definition of a strong class, if $B\leq C\in\mathcal{K}$, then $\binom{B}{A}=\binom{C}{A}\cap\mathcal{P}(B)$.)
For $A,B\in\mathcal{K}$, $r\in\mathbb{N}_{+}$, an $r$-colouring of the set $\binom{B}{A}$ is a function $\chi:\binom{B}{A}\to\{1,\cdots,r\}$. We say that $\binom{B}{A}$ is monochromatic in the $r$-colouring $\chi$ if $\chi$ is constant on $\binom{B}{A}$.
For $A,B,C\in\mathcal{K}$, $r\in\mathbb{N}_{+}$, we write $C\rightarrow(B)^{A}_{r}$ if for every $r$-colouring $\chi$ of $\binom{C}{A}$, there exists $B^{\prime}\in\binom{C}{B}$ such that $\binom{B^{\prime}}{A}$ is monochromatic in $\chi$.
We say that $(\mathcal{K},\leq)$ has the Ramsey property if for $r\in\mathbb{N}_{+}$, $A,B\in\mathcal{K}$, there exists $C\in\mathcal{K}$ with $C\rightarrow(B)^{A}_{r}$.
The following is a well-known observation of Nešetřil ([16]) which shows a strong connection between Fraïssé theory and structural Ramsey theory (see [4, Theorem 2.11] for a proof):
Proposition 2.19.
Let $(\mathcal{K},\leq)$ be a strong, rigid class with JEP and the Ramsey property. Then $(\mathcal{K},\leq)$ has the amalgamation property.
The following is one of the key results of [13], and establishes a link between topological dynamics and structural Ramsey theory:
Theorem 2.20 ([13], Th. 4.8; [17], Th. 1).
Let $(\mathcal{K},\leq)$ be an amalgamation class of rigid $L$-structures, with Fraïssé limit $M$ and $G=\operatorname{Aut}(M)$. Then $G$ is extremely amenable iff $(\mathcal{K},\leq)$ has the Ramsey property.
Here, a finite $L$-structure is rigid if it has trivial automorphism group.
2.8. Reasonable expansions and the expansion property
In this section, $L^{+}$ is an expansion of the language $L$.
Here we follow Section 2.3 of [4], which takes the notion of a reasonable expansion from [19].
Definition 2.21.
Let $(\mathcal{K},\leq)$ be an amalgamation class of $L$-structures. A class $\mathcal{D}$ of finite $L^{+}$-structures is a reasonable expansion of $(\mathcal{K},\leq)$ if $\mathcal{D}$ is closed under isomorphisms and satisfies the following:
  1. (1)
    $\mathcal{K}$ is the class of $L$-reducts of $\mathcal{D}$;
  2. (2)
    for $A\in\mathcal{K}$, $A$ has finitely many expansions in $\mathcal{D}$ (weak coprecompactness);
  3. (3)
    for $B^{+}\in\mathcal{D}$, if $A^{+}\subseteq B^{+}$ and $A^{+}|_{L}\leq B^{+}|_{L}$, then $A^{+}\in\mathcal{D}$;
  4. (4)
    for $f:A\to B$ a strong embedding in $(\mathcal{K},\leq)$, if $A^{+}\in\mathcal{D}$ is an expansion of $A$, then there exists an expansion $B^{+}\in\mathcal{D}$ of $B$ such that $f:A^{+}\to B^{+}$ is an embedding.
Note that if $(\mathcal{K}^{+},\leq^{+})$ is a strong expansion of $(\mathcal{K},\leq)$, then $\mathcal{K}^{+}$ satisfies properties (1) and (4) above.
Lemma 2.22.
For any amalgamation class $(\mathcal{K},\leq)$ of $L$-structures, the order expansion $\mathcal{K}^{\prec}$ of $(\mathcal{K},\leq)$ (see Definition 2.12) is a reasonable expansion of the class $(\mathcal{K},\leq)$.
The proof is straightforward.
Lemma 2.23.
The class $\mathcal{D}_{1}$ is a reasonable expansion of $(\mathcal{C}_{1},\leq_{1})$.
Proof.
As $(\mathcal{D}_{1},\leq_{s})$ is a strong expansion of $(\mathcal{C}_{1},\leq_{1})$, we have that $\mathcal{D}_{1}$ satisfies (1) and (4) in the definition of reasonableness. Parts (2) and (3) are immediate. ∎
Definition 2.24.
Let $(\mathcal{K},\leq)$ be an amalgamation class with Fraïssé limit $M$. Let $\mathcal{D}$ be a reasonable $L^{+}$-expansion of $(\mathcal{K},\leq)$.
Let $X(\mathcal{D})$ be the set consisting of the $L^{+}$-expansions $M^{+}$ of $M$ such that $M^{+}|_{A}\in\mathcal{D}$ for all $A\leq M$. (Here, $M^{+}|_{A}$ denotes the $L^{+}$-structure induced on the domain of $A$ by $M^{+}$.) We define a topology on $X(\mathcal{D})$: for $B\leq M$ with expansion $B^{+}\in\mathcal{D}$, we specify a basic open set $U(B^{+})=\{M^{+}\in X(\mathcal{D}):M^{+}|_{B}=B^{+}\}$. We will see below that $X(\mathcal{D})$ gives a $G$-flow.
Lemma 2.25.
Let $(\mathcal{K},\leq)$ be an amalgamation class with Fraïssé limit $M$. Let $\mathcal{D}$ be a reasonable $L^{+}$-expansion of $(\mathcal{K},\leq)$. Then for $A\leq M$ with expansion $A^{+}\in\mathcal{D}$, there is an expansion $M^{+}\in X(\mathcal{D})$ with $M^{+}|_{A}=A^{+}$.
Proof.
Write $M$ as the union of an increasing $\leq$-chain $A=A_{1}\leq A_{2}\leq\cdots$ starting at $A$. Let $A^{+}_{1}=A^{+}$. By part (4) of reasonableness, we may inductively expand $A_{i}$ to some $A^{+}_{i}$ ($i\geq 2$) such that $A^{+}_{i-1}\leq A^{+}_{i}\in\mathcal{D}$. We then take $M^{+}=\bigcup_{i\geq 1}A^{+}_{i}$. For $B\leq M$, we have that $B\leq A_{i}$ for some $i$, so by part (3) of reasonableness, $M^{+}|_{B}\in\mathcal{D}$. So $M^{+}\in X(\mathcal{D})$. ∎
Lemma 2.26 (after Prop. 5.3, [19]).
Let $\mathcal{D}$ be a reasonable expansion of the amalgamation class $(\mathcal{K},\leq)$. Let $M$ be the Fraïssé limit of $(\mathcal{K},\leq)$ and let $G=\operatorname{Aut}(M)$. Then $X(\mathcal{D})$ is a $G$-flow with the natural action.
Proof.
By Lemma 2.25, $X(\mathcal{D})$ is non-empty. It is straightforward to see that the action is continuous and $X(\mathcal{D})$ is Hausdorff. To see that $X(\mathcal{D})$ is compact, consider the topological space $Q=\prod_{A\leq M}\{A^{+}\mid A^{+}\text{ is an expansion of }A\,\wedge\,A^{+}\in\mathcal{D}\}$, where each set of the product has the discrete topology and $Q$ has the product topology. Using part (2) of reasonableness, we have that $Q$ is compact. We define a map $\gamma:X(\mathcal{D})\to Q$, $\gamma(M^{+})=(M^{+}|_{A})_{A\leq M}$, and it is straightforward to see that $\gamma$ is a homeomorphism onto its image. Thus it suffices to show that $\operatorname{Im}(\gamma)$ is a closed subspace of $Q$. We have that
$\operatorname{Im}(\gamma)=\{(p_{A})_{A\leq M}\in Q:\forall\,B\leq C\leq M,\,p_{C}|_{B}=p_{B}\},$
and so if an element of $Q$ lies in the complement of $\operatorname{Im}(\gamma)$, this is witnessed on a finite set. ∎
Definition 2.27.
Let $(\mathcal{K},\leq)$ be an amalgamation class with Fraïssé limit $M$ and $G=\operatorname{Aut}(M)$. As the order expansion $\mathcal{K}^{\prec}$ is a reasonable expansion of $(\mathcal{K},\leq)$, we have that $X(\mathcal{K}^{\prec})$ is a $G$-flow: we denote this by $\mathcal{LO}(M)$, the flow of linear orders on $M$.
As $\mathcal{D}_{1}$ is a reasonable expansion of $(\mathcal{C}_{1},\leq_{1})$, we have that $X(\mathcal{D}_{1})$ is a $G_{1}$-flow, which we denote by $\operatorname{Or}(M_{1})$, the flow of orientations. (Recall that $M_{1}$ is the Fraïssé limit of $(\mathcal{C}_{1},\leq_{1})$ and $G_{1}=\operatorname{Aut}(M_{1})$.)
Lemma 2.28.
Let $\mathcal{D}$ be a reasonable expansion of the amalgamation class $(\mathcal{K},\leq)$. Let $M$ be the Fraïssé limit of $(\mathcal{K},\leq)$ and let $G=\operatorname{Aut}(M)$.
Let $Y$ be a subflow of $X(\mathcal{D})$. Let $\mathcal{D}^{\prime}\subseteq\mathcal{D}$ be the class of finite $L^{+}$-structures which $\leq$-embed into some element of $Y$. Then
  1. (1)
    $\mathcal{D}^{\prime}$ is a reasonable expansion of $(\mathcal{K},\leq)$ with $X(\mathcal{D}^{\prime})=Y$;
  2. (2)
    if $Y=\overline{G\cdot M_{0}^{+}}$ for some $M_{0}^{+}\in X(\mathcal{D})$, then $\mathcal{D}^{\prime}=\operatorname{Age}_{\leq}(M_{0}^{+})$.
(Part (1) is from [4, Lemma 2.16].)
Proof.
(1): Parts (1), (2), (3) of reasonableness of $\mathcal{D}^{\prime}$ are straightforward, and (4) follows from the $\leq$-homogeneity of $M$ and the $G$-invariance of $Y$. Clearly $Y\subseteq X(\mathcal{D}^{\prime})$. To see that $X(\mathcal{D}^{\prime})\subseteq Y$, take $M^{+}\in X(\mathcal{D}^{\prime})$. It remains to show that $M^{+}\in\operatorname{cl}_{X(\mathcal{D}^{\prime})}(Y)=Y$: this follows from the $\leq$-homogeneity of $M$ and the $G$-invariance of $Y$.
(2): Let $A^{+}\in\mathcal{D}^{\prime}$. By the definition of $\mathcal{D}^{\prime}$ we may assume $A^{+}\leq M_{1}^{+}$ for some $M_{1}^{+}\in Y$. As $Y=\overline{G\cdot M_{0}^{+}}$, there is $g\in G$ such that $A^{+}=M_{1}^{+}|_{A}=(gM_{0}^{+})|_{A}$, so $A^{+}$leq-embeds into $M_{0}^{+}$ and thus $\mathcal{D}^{\prime}\subseteq\operatorname{Age}_{\leq}(M_{0}^{+})$. The reverse inclusion is immediate as $M_{0}^{+}\in Y$. ∎
Definition 2.29.
Let $\mathcal{D}$ be a reasonable expansion of the amalgamation class $(\mathcal{K},\leq)$. We say that $\mathcal{D}$ has the expansion property over $(\mathcal{K},\leq)$ if for $A\in\mathcal{K}$, there exists $B$ in $\mathcal{K}$ with $A\leq B$ such that for all expansions $A^{+},B^{+}$ of $A,B$ in $\mathcal{D}$, there exists a $\leq$-embedding $A^{+}\to B^{+}$.
It is easy to see that to prove the expansion property for $\mathcal{D}$, it suffices to show that for $A^{+}\in\mathcal{D}$, there exists $B\in\mathcal{K}$ such that for all expansions $B^{+}\in\mathcal{D}$ of $B$, there exists a $\leq$-embedding $A^{+}\to B^{+}$ (use weak coprecompactness of $\mathcal{D}$ and JEP for $\mathcal{K}$).
Theorem 2.30 ([17], Th. 4).
Let $\mathcal{D}$ be a reasonable expansion of the amalgamation class $(\mathcal{K},\leq)$ with Fraïssé limit $M$ and $G=\operatorname{Aut}(M)$.
Then the $G$-flow $X(\mathcal{D})$ is minimal iff $\mathcal{D}$ has the expansion property over $(\mathcal{K},\leq)$.
2.9. Meagre orbits
We now introduce the weak amalgamation property (WAP), which will be essential in the proof of Theorem 4.1. (WAP was first defined in [14]. A similar property (the almost amalgamation property) was previously defined in [12].)
Definition 2.31.
Let $\mathcal{D}$ be a reasonable class of $L^{+}$-expansions of an $L$-amalgamation class $(\mathcal{K},\leq)$. We say that $(\mathcal{D},\leq)$ has the weak amalgamation property (WAP) if:
for all $A\in\mathcal{D}$, there exists $B\in\mathcal{D}$ and a $\leq$-strong $L^{+}$-embedding $f:A\to B$ such that, for any $\leq$-strong $L^{+}$-embeddings $f_{i}:B\to C_{i}\in\mathcal{D}$ ($i=0,1$), there exists $D\in\mathcal{D}$ and $\leq$-strong $L^{+}$-embeddings $g_{i}:C_{i}\to D$ ($i=0,1$) with $g_{0}\circ f_{0}\circ f=g_{1}\circ f_{1}\circ f$. (Note that here we specify only that the diagram commutes for $A$.)
[…]
Lemma 2.32 ([4, Lemma 2.23] - adapting [14, Theorem 3.4]).
Let $\mathcal{D}$ be a reasonable class of $L^{+}$-expansions of an $L$-amalgamation class $(\mathcal{K},\leq)$. Let $M$ be the Fraïssé limit of $(\mathcal{K},\leq)$ and let $G=\operatorname{Aut}(M)$. Suppose that $X(\mathcal{D})$ is a minimal flow.
If $(\mathcal{D},\leq)$ does not have the weak amalgamation property, then all $G$-orbits on $X(\mathcal{D})$ are meagre.
For a proof, see [18, Lemma 1.77].
3. Admissible orders: a Ramsey expansion of $(\mathcal{D}_{1},\sqsubseteq_{s})$
We now provide an explicit description of a Ramsey expansion of $(\mathcal{D}_{1},\sqsubseteq_{s})$, given by the admissible orders on $(\mathcal{D}_{1},\sqsubseteq_{s})$, using Theorem 1.4 of [5]. This will be an essential tool in the proof of 4.1. (This Ramsey expansion will also have the expansion property over $(\mathcal{D}_{1},\sqsubseteq_{s})$, though we will not use this in the proof of Theorem 4.1.) In the below two definitions, we adapt definitions from [5, Section 1 and Section 3] to the specific case of $(\mathcal{D}_{1},\sqsubseteq_{s})$.
Definition 3.1.
Let $A\in\mathcal{D}_{1}$. For $a\in A$, let $a^{\circ}=\operatorname{scl}_{A}(a)\setminus\{a\}$.
For $a\in A$, we inductively define the level $\operatorname{l}_{A}(a)$ of $a$ as follows. If $\operatorname{scl}_{A}(a)=\{a\}$, then $\operatorname{l}_{A}(a)=0$. Otherwise, let $b$ be a vertex of $a^{\circ}$ of maximum level, and then define $\operatorname{l}_{A}(a)=\operatorname{l}_{A}(b)+1$. We write $\operatorname{L}_{n}(A)$ ($n\geq 0$) for the set of vertices of $A$ of level $n$.
We say that $a,b\in A$ are homologous if $a^{\circ}=b^{\circ}$ and there is an isomorphism $\operatorname{scl}_{A}(a)\to\operatorname{scl}_{A}(b)$ which is the identity on $a^{\circ}=b^{\circ}$. We let $Q_{A}(a)$ denote the set of vertices of $A$ homologous to $a$, and call $Q_{A}(a)$ the cone of $a$.
If there is $a\in A$ with $A=\operatorname{scl}_{A}(a)$, we call $A$ a closure-extension with head vertex $a$, and write $A^{\circ}=a^{\circ}$. (Note that $a$ is necessarily unique.)
Definition 3.2.
[…]
  • ($\ast$)
    if $|A|<|B|$, then $A^{\prec}\lhd B^{\prec}$.
We say that a class $\mathcal{O}\subseteq\mathcal{D}_{1}^{\prec}$ is a class of admissible orderings of structures in $\mathcal{D}_{1}$ if:
  1. (1)
    each $A\in\mathcal{D}_{1}$ has an expansion $A^{\prec}\in\mathcal{O}$;
  2. (2)
    $\mathcal{O}$ is closed under $\sqsubseteq_{s}$-substructures;
  3. (3)
    for $A^{\prec}\in\mathcal{O}$ and $u,v\in A$, if:
    • $\operatorname{scl}_{A}(u)^{\prec}\lhd\operatorname{scl}_{A}(v)^{\prec}$, or
    • $\operatorname{scl}_{A}(u)^{\prec}\cong\operatorname{scl}_{A}(v)^{\prec}$ and $u^{\circ}$ is lexicographically before $v^{\circ}$ in the order $\prec_{A}$,
    then $u\prec_{A}v$;
  4. (4)
    for each $B\in\mathcal{D}_{1}$, if $A_{1},\cdots,A_{n}\sqsubseteq_{s}B$ and $\prec^{\prime}$ is a linear order on $A=\bigcup_{i\leq n}A_{i}$ such that $\prec^{\prime}$ satisfies (3) and each $A_{i}$ is admissibly ordered by $\prec^{\prime}$, then there exists an admissible order $\prec_{B}$ on $B$ extending $\prec^{\prime}$;
(In the above, we adapt [5, Definition 3.5]. Several aspects of the general definition in [5] simplify in this case: (A4) can be omitted as closure components are single vertices, and (A6) follows from ()($\ast$)( ∗ ) and $(3)$.)
The below theorem is an immediate translation of [5, Theorem 1.4] to the context of this paper. We will explain how to adapt [5, Theorem 1.4] to our context at the end of this section.
Proposition 3.3.
There exists a class $\mathcal{O}_{1}\subseteq\mathcal{D}_{1}^{\prec}$ of admissible orderings of structures in $\mathcal{D}_{1}$. We have that $(\mathcal{O}_{1},\sqsubseteq_{s})$ is an amalgamation class, and $(\mathcal{O}_{1},\sqsubseteq_{s})$ has the Ramsey property and the expansion property over $(\mathcal{D}_{1},\sqsubseteq_{s})$.
Lemma 3.4.
$(\mathcal{O}_{1},\sqsubseteq_{s})$ is a strong expansion of $(\mathcal{D}_{1},\sqsubseteq_{s})$.
Proof.
Parts (1) and (2) in the definition of strong expansion are immediate. Part (3) follows immediately from part (4) of the definition of admissible orders. ∎
We now give a specific property resulting from 3.2 that we will use in the proof of 4.1, the main result of this paper.
Lemma 3.5.
Let $A^{\prec}\in\mathcal{O}_{1}$. Let $a\in A$ and let $b\in a^{\circ}$. Then $b\prec_{A}a$.
Proof.
As $|\operatorname{scl}_{A}(b)|<|\operatorname{scl}_{A}(a)|$, by parts ()($\ast$)( ∗ ) and $(3)$ in 3.2 we have $b\prec_{A}a$. ∎
We now explain how to adapt [5, Theorem 1.4] and the definition of admissible orders found in [5] to give the definitions and theorem above. The paper [5] gives Ramsey expansions for classes of finite structures in languages that may include set-valued function symbols, which enables us to deal with the strong class $(\mathcal{D}_{1},\sqsubseteq_{s})$.
Definition 3.6 ([5, Section 1]).
A language $L=L_{R}\cup L_{F}$ of relation and set-valued function symbols consists of a set $L_{R}$ of relation symbols and a set $L_{F}$ of set-valued function symbols $L_{F}$, where each symbol has an associated arity $n\in\mathbb{N}_{+}$.
An $L$-structure $(A,(R_{A})_{R\in L_{R}},(F_{A})_{F\in L_{F}})$ consists of a set $A$ (the domain) together with sets $R_{A}\subseteq A^{n}$ for each relation symbol $R\in L_{R}$ of arity $n$ and functions $F_{A}:A^{n}\to\mathcal{P}(A)$ for each set-valued function symbol $F\in L_{F}$ of arity $n$. Usually we will just write $A$ to denote the structure.
A function $f:A\to B$ between $L$-structures $A,B$ is an embedding if $f$ is injective and:
  • for each relation symbol $R\in L_{R}$ of arity $n$,
    $(a_{1},\cdots,a_{n})\in R_{A}\Leftrightarrow(f(a_{1}),\cdots,f(a_{n}))\in R_{B};$
  • for each set-valued function symbol $F\in L_{F}$ of arity $n$,
    $f(F_{A}(a_{1},\cdots,a_{n}))=F_{B}(f(a_{1}),\cdots,f(a_{n})).$
For $L$-structures $A,B$, we say that $A$ is a substructure of $B$, written $A\subseteq B$, if the domain of $A$ is a subset of the domain of $B$ and the inclusion map $A\hookrightarrow B$ is an embedding of $L$-structures.
We define the hereditary property, joint embedding property, amalgamation property, Ramsey property and expansion property for classes of $L$-structures exactly as for usual first-order languages, and we also define amalgamation classes and Ramsey classes as before.
Let $A\subseteq B_{0},B_{1}$ be $L$-structures, and suppose that $B_{0}\cap B_{1}=A$. The free amalgam of $B_{0},B_{1}$ over $A$ is the $L$-structure $C$ with domain $B_{0}\cup B_{1}$, where $R_{C}=R_{B_{0}}\cup R_{B_{1}}$ for each $R\in L_{R}$ and where, for each $F\in L_{F}$ of arity $n$, the function $F_{C}:C^{n}\to\mathcal{P}(C)$ is defined by $F_{C}(\bar{c})=F_{B_{i}}(\bar{c})$ for $\bar{c}\in B_{i}^{n}$ ($i=0,1$) and $F_{C}(\bar{c})=\varnothing$ otherwise. An amalgamation class where amalgams can always be taken to be free amalgams is called a free amalgamation class.
The above framework enables us to deal with $(\mathcal{D}_{1},\sqsubseteq_{s})$ as follows ([5, Section 5.1]). Let $\tilde{L}$ consist of the relational language $L_{\text{or}}$ of oriented graphs together with a unary set-valued function symbol $F$. Let $\tilde{\mathcal{D}}_{1}$ consist of the $\tilde{L}$ structures $\tilde{A}=(A,F_{A})$ where $A\in\mathcal{D}_{1}$ and $F_{A}:A\to\mathcal{P}(A)$ is a function sending each vertex of $A$ to its out-neighbourhood in $A$. Then there is a bijection $\mathcal{D}_{1}\to\tilde{\mathcal{D}}_{1}$ sending each $A\in\mathcal{D}_{1}$ to its unique $\tilde{L}$-expansion $\tilde{A}$ in $\tilde{\mathcal{D}}_{1}$, and for $A,B\in\mathcal{D}_{1}$, we have that $A\sqsubseteq_{s}B$ iff $\tilde{A}\subseteq\tilde{B}$. We then have that $\tilde{L}$-embeddings between elements of $\tilde{\mathcal{D}}_{1}$ are $\sqsubseteq_{s}$-embeddings when considered in the language $L_{\text{or}}$, and therefore $\tilde{\mathcal{D}}_{1}$ is a free amalgamation class.
We now recall [5, Theorem 1.4], which will give us an explicitly defined Ramsey expansion of $\tilde{\mathcal{D}}_{1}$ via admissibly ordered structures.
Theorem ([5, Theorem 1.4]).
Let $L$ be a language (consisting of relation and set-valued function symbols). Let $\mathcal{K}$ be a free amalgamation class of $L$-structures. Then there exists an explicitly defined amalgamation class $\mathcal{O}\subseteq\mathcal{K}^{\prec}$ of admissible orderings such that:
  • every $A\in\mathcal{K}$ has an ordering in $\mathcal{O}$;
  • the class $\mathcal{O}$ has the Ramsey property and the expansion property over $\mathcal{K}$.
The above theorem, together with [5, Definition 3.5], which gives the explicit definition of admissible orders, gives a Ramsey expansion $\tilde{\mathcal{O}}_{1}$ of $\tilde{\mathcal{D}}_{1}$. Using the correspondence between $\tilde{\mathcal{D}}_{1}$ and $(\mathcal{D}_{1},\sqsubseteq_{s})$ detailed in the preceding paragraph, we thus obtain a Ramsey expansion $(\mathcal{O}_{1},\sqsubseteq_{s})$ of $(\mathcal{D}_{1},\sqsubseteq_{s})$ satisfying the conditions of 3.2 (this definition is just a direct adaptation of [5, Definition 3.5]).
4. Minimal subflows of $M_{1}$
In this section, we prove the main theorem of this paper:
Theorem 4.1.
Let $Y\subseteq\mathcal{LO}(M_{1})$ be a minimal subflow of $\mathcal{LO}(M_{1})$. Then all $G$-orbits on $Y$ are meagre.
4.1. Preparatory definitions and lemmas
Definition 4.2.
Let $N_{1}=(M_{1},\rho)$ be the Fraïssé limit of $(\mathcal{D}_{1},\sqsubseteq_{s})$, and let $(N_{1},\prec_{\alpha})$ be the Fraïssé limit of $(\mathcal{O}_{1},\sqsubseteq_{s})$. (Here, we use 2.11: $(\mathcal{D}_{1},\sqsubseteq_{s})$ is a strong expansion of $(\mathcal{C}_{1},\leq_{1})$ and $(\mathcal{O}_{1},\sqsubseteq_{s})$ is a strong expansion of $(\mathcal{D}_{1},\sqsubseteq_{s})$.)
Recall that $G_{1}=\operatorname{Aut}(M_{1})$. Let $H_{1}=\operatorname{Aut}(N_{1},\alpha)$. By 3.3 and 2.20, we have that $H_{1}$ is extremely amenable.
We will write $G=G_{1},H=H_{1}$ in the remainder of Section 4 for ease of notation.
Definition 4.3.
Let $a\in N_{1}$. As $N_{1}$ is the union of an increasing chain of $\sqsubseteq_{s}$-substructures, we have that $\operatorname{scl}_{N_{1}}(a)$ is finite, and for any $A\sqsubseteq_{s}N_{1}$ we have $\operatorname{scl}_{N_{1}}(a)=\operatorname{scl}_{A}(a)$. We define $a^{\circ}=\operatorname{scl}_{N_{1}}(a)\setminus\{a\}$, and define homologous vertices and cones in $N_{1}$ as in 3.1. We define the level $\operatorname{l}_{N_{1}}(a)$ of $a$ in $N_{1}$, usually just denoted $\operatorname{l}(a)$, to be the level of $a$ in $\operatorname{scl}_{N_{1}}(a)$.
Lemma 4.4.
Let $\prec_{\beta}$ be an $H$-fixed point in the flow $H\curvearrowright\mathcal{LO}(M_{1})$, and let $Q$ be a cone of $N_{1}$. Then $\prec_{\beta}$ agrees with either $\prec_{\alpha}$ or $\prec_{\alpha}^{\prime}$ on $Q$, where $\prec_{\alpha}^{\prime}$ denotes the reverse of the linear order $\prec_{\alpha}$.
Proof.
Take $a_{0},b_{0}\in Q$ with $a_{0}\prec_{\alpha}b_{0}$. Then for $a,b\in Q$ with $a\prec_{\alpha}b$, by 3.5 there exists an ordered digraph isomorphism $f:\operatorname{scl}_{N_{1}}(a_{0},b_{0})^{\prec_{\alpha}}\to\operatorname{scl}_{N_{1}}(a,b)^{\prec_{\alpha}}$ with $f(a_{0})=a,f(b_{0})=b$, and by $\sqsubseteq_{s}$-ultrahomogeneity we may extend to an element $f\in H$.
As $H\subseteq G_{\beta}$, $f$ is $\beta$-preserving. If $a_{0}\prec_{\beta}b_{0}$, then $f(a_{0})\prec_{\beta}f(b_{0})$, so $a\prec_{\beta}b$, and so $\prec_{\beta}$ agrees with $\prec_{\alpha}$ on $Q$. If $a_{0}\succ_{\beta}b_{0}$, then $\prec_{\beta}$ agrees with $\prec_{\alpha}^{\prime}$ on $Q$. ∎
4.2. Setup and proof notation
Before beginning the proof, we first need to set up our approach.
Let $Y\subseteq\mathcal{LO}(M_{1})$ be a minimal subflow of $G\curvearrowright\mathcal{LO}(M_{1})$. As $H$ is extremely amenable, the flow $H\curvearrowright Y$ has an $H$-fixed point $\prec_{\beta}$, and as $Y$ is a minimal $G$-flow, we have $Y=\overline{G\,\cdot\!\prec_{\beta}}$. Let $\mathcal{J}=\operatorname{Age}_{\leq_{1}}(M_{1},\prec_{\beta})$. By 2.28, we have $Y=X(\mathcal{J})$. We will show that $(\mathcal{J},\leq_{1})$ does not have the weak amalgamation property (WAP), which implies that all $G$-orbits on $Y$ are meagre by Lemma 2.32.
We will now use the above notation throughout the rest of this section.
4.3. Proof idea - informal overview
We will assume $(\mathcal{J},\leq_{1})$ has WAP, for a contradiction. Let $\{a_{0}\}\in\mathcal{J}$ be a singleton with the trivial linear order. By assumption $\{a_{0}\}$ has a WAP-witness $A^{\prec}$. We will then construct $\leq_{1}$-embeddings of $A^{\prec}$ into two ordered graphs $C_{0}^{\prec},C_{1}^{\prec}\in\mathcal{J}$ which are WAP-incompatible: it will not be possible to find $D^{\prec}$ completing the WAP commutative diagram for $\{a_{0}\}$ with the two embeddings, and this will give a contradiction.
The incompatibility of the two ordered graphs $C_{0}^{\prec},C_{1}^{\prec}$ in $\mathcal{J}$ will result from them forcing incompatible orientations: we can use the order $\prec_{\beta}$ to force certain edge orientations in $\rho$. The incompatible orientations will consist of a binary out-directed tree $T_{0}$ and a binary out-directed tree with the successor-closures of two vertices identified, which we denote by $T_{1}$: these cannot start from the same point of a $2$-orientation, as one contains a $4$-cycle and the other does not. The idea to use two incompatible orientations in the WAP commutative diagram comes from the proof of [4, Theorem 5.2].
The key difficulties in the proof of Theorem 4.1 are showing that we can use $\prec_{\beta}$ (specifically, particular finite ordered graphs in $\mathcal{J}=\operatorname{Age}_{\leq_{1}}(M_{1},\prec_{\beta})$) to force orientations of edges in $\rho$ (Lemma 4.6), and also showing that the ordered graphs that we construct to force orientations of edges in $\rho$ do in fact lie in $\mathcal{J}$ (Lemma 4.7).
4.4. Attaching trees and near-trees
For $q\in\mathbb{N}_{+}$, let $T_{0}(q)$ be the digraph given by a binary tree of height $12q+$ 1, oriented outwards towards the leaves and with head vertex $c$. Let $T_{1}(q)$ be the digraph given by taking $T_{0}(q)$ and identifying the successor-closures of two vertices at height $q+2$ whose paths to the head vertex $c$ meet at height $q$. We have $T_{0}(q),T_{1}(q)\in\mathcal{D}_{1}$.
Let $T$ be one of the digraphs $T_{0}(q)$ or $T_{1}(q)$. Take $C\in\mathcal{D}_{1}$ with each vertex having out-degree $2$ or 00. Let $D_{T}$ be the digraph consisting of $C$ together with, for each vertex $v\in C$ with $\operatorname{d_{{}_{+}}\!}(v)=0$, a copy of $T$ attached at $v$, where we identify $c$ and $v$. Let $Z_{T}$ denote the sub-digraph of $D_{T}$ whose vertices are the vertices of the copies of $T$ attached to $C$ in $D_{T}$. Let $D_{T}^{-}$ denote the graph reduct of $D_{T}$. (We will use this notation throughout this section. See Figure 1.)
We have $D_{T}\in\mathcal{D}_{1}$. Let ${D_{T}}^{\prime}$ be the acyclic $2$-reorientation of $D_{T}$ where the copies of $T$ have been oriented so that the non-head vertices of each copy of $T$ are directed towards the head vertex $c$, leaving the orientation on vertices of $C$ unchanged. Then we have $C\sqsubseteq_{s}{D_{T}}^{\prime}$ in this reorientation, and so $C^{-}\leq_{1}D_{T}^{-}$.
Definition 4.5.
Let $C\in\mathcal{D}_{1}$ with each vertex having out-degree $2$ or 00, and let $D_{T}$ be defined as above.
An ordered graph $C_{T}^{\prec}\in\mathcal{J}$ is a $T$-witness ordered graph for C if:
  • $C_{T}$ consists of the graph reduct ${D_{T}}^{-}$ of $D_{T}$ together with, for each non-leaf tree vertex $v$ of $D_{T}$, an additional $10$ copies of $\operatorname{scl}_{D_{T}}(v)$ freely amalgamated (as graphs) over $\operatorname{scl}_{D_{T}}(v)^{\circ}$, and $C\leq_{1}C_{T}$;
  • for each non-leaf tree vertex $v\in D_{T}$, the additional $10$ copies of $v$ may be labelled as $v_{-5},\cdots,v_{-1},v_{1},\cdots,v_{5}$ so that $v_{-5}\prec\cdots\prec v_{-1}\prec v\prec v_{1}\prec\cdots\prec v_{5}$ in $\prec_{C_{T}}$. (We call these $v_{i}$ the witness vertices of $v$.)
(See Figure 2.)
[…]
Figure 1. The oriented graph $D_{T}$.
[…]
Figure 2. The ordered graph $C_{T}^{\prec}$ with witness vertices indicated on one vertex.
The following is the key lemma here.
Lemma 4.6.
Let $C\in\mathcal{D}_{1}$ with each vertex having out-degree $2$ or 00, and let $C_{T}^{\prec}\in\mathcal{J}$ be a $T$-witness ordered graph for $C$. As $C_{T}^{\prec}\in\mathcal{J}$, there exists a $\leq_{1}$-ordered graph embedding $\theta:C_{T}^{\prec}\to(M_{1},\prec_{\beta})$. Then, considering the digraph structure on $Z_{T}$ induced by $D_{T}$, $\theta|_{Z_{T}}:Z_{T}\to(M_{1},\rho)$ is also a digraph embedding.
Proof.
We may take $\theta=\operatorname{id}$ for ease of notation. Take $v,x,y\in Z_{T}$ with out-edges $vx,vy$ in the orientation of $Z_{T}$. We need to show that $v$ has out-edges $vx,vy$ in the orientation $\rho$ of $M_{1}$. Let $v_{-5},\cdots,v_{-1},v_{1},\cdots,v_{5}$ be the witness vertices of $v$ in $C_{T}^{\prec}$, and let $v_{0}=v$. As $\theta$ is a $\leq_{1}$-ordered graph embedding, we have that $v_{i}\prec_{\beta}v_{j}$ for $i<j$, and we have undirected edges $v_{i}x,v_{i}y$ for $-5\leq i\leq 5$.
As $\rho$ is a $2$-orientation, for some $i$ with $-5\leq i\leq-1$ we must have that $v_{i}x,v_{i}y$ are out-edges of $\rho$, and likewise for some $j$ with $1\leq j\leq 5$ we must have that $v_{j}x,v_{j}y$ are out-edges of $\rho$. If either $xv_{0}\in\rho$ or $yv_{0}\in\rho$, then $v_{0}\in\operatorname{scl}_{\rho}(x,y)$, and as $v_{i},v_{j}$ lie in the same cone, by 3.5 there exists $h\in H$ with $hv_{i}=v_{j}$ and $h$ fixing $v_{0}$. As $H\subseteq G_{\beta}$, we have that $h\in G_{\beta}$. But $v_{i}\prec_{\beta}v_{0}$, so $hv_{i}\prec_{\beta}hv_{0}$, thus $v_{j}\prec_{\beta}v_{0}$ contradiction. So therefore $v_{0}x,v_{0}y\in\rho$. ∎
Lemma 4.7.
Let $C\in\mathcal{D}_{1}$ with each vertex having out-degree $2$ or 00. Then there exists a $T$-witness ordered graph $C_{T}^{\prec}\in\mathcal{J}$ for $C$.
Proof.
Let $d_{1},\cdots,d_{k}$ be an enumeration of the non-leaf tree vertices of $D_{T}$ which preserves the order of levels, i.e. for $i<j$, $\operatorname{l}_{D_{T}}(d_{i})\leq\operatorname{l}_{D_{T}}(d_{j})$. We will show, by induction on $i$, that for $0\leq i\leq k$ there exists an ordered graph $C_{i}^{\prec}\in\mathcal{J}$ such that:
  1. (1)
    $C_{i}$ consists of $D_{T}$ together with, for $1\leq j\leq i$, an additional $10$ copies of $\operatorname{scl}_{D_{T}}(d_{j})$ freely amalgamated (as graphs) over $\operatorname{scl}_{D_{T}}(d_{j})^{\circ}$;
  2. (2)
    for $1\leq j\leq i$, the $10$ copies of $d_{j}$ may be labelled as $d_{j,-5},\cdots,d_{j,-1},d_{j,1},\cdots,d_{j,5}$ such that $d_{j,-5}\prec\cdots\prec d_{j,-1}\prec d_{j}\prec d_{j,1}\prec\cdots\prec d_{j,5}$ in $\prec_{C_{i}}$. We will call these the witness vertices of $d_{j}$, and let $W_{j}$ denote the set of witness vertices of $d_{j}$.
For the base case $i=0$, take $C_{0}={D_{T}}^{-}$. As $C_{0}\in\mathcal{C}_{1}$ and $\mathcal{J}$ is a reasonable class of expansions of $(\mathcal{C}_{1},\leq_{1})$, there exists a linear order $\prec_{C_{0}}$ on $C_{0}$ such that $C_{0}^{\prec}\in\mathcal{J}$, and then $C_{0}^{\prec}$ satisfies (1) and (2) vacuously.
For the induction step, assume we have $C_{i}^{\prec}\in\mathcal{J}$ satisfying (1) and (2). Let
$X=\operatorname{L}_{0}(D_{T})\cup\bigcup_{1\leq j\leq i}\operatorname{scl}_{D_{T}}(d_{j})\cup\bigcup_{1\leq j\leq i}W_{j}.$
There is an acyclic $2$-orientation $\tau_{i}$ of $C_{i}$ in which $X$ is successor-closed: take the orientation of $D_{T}$, and orient the two edges of each witness vertex $d_{j,m}$ outwards from $d_{j,m}$. Thus $X\leq_{1}C_{i}$. Note that for $j^{\prime}>i\geq j$ we have $\operatorname{l}_{D_{T}}(d_{j^{\prime}})\geq\operatorname{l}_{D_{T}}(d_{j})$, so $d_{j^{\prime}}\notin X$ for $j^{\prime}>i$.
Let $(E,\tau)$ be the free amalgam of $(C_{i},\tau_{i})$ $11$ times over $(X,\tau_{i})$. As $\mathcal{D}_{1}$ is a free amalgamation class, we have $(E,\tau)\in\mathcal{D}_{1}$. Hence $E\in\mathcal{C}_{1}$, and we have $X\leq E$.
Let $\prec_{X}\,=\,\prec_{C_{i}}\!|_{X}$. We have that $X^{\prec}\in\mathcal{J}$, so let $\theta_{X}:X^{\prec}\to(M_{1},\prec_{\beta})$ be a $\leq_{1}$-ordered graph embedding. By the extension property of $M_{1}$, we have a $\leq_{1}$-graph embedding $\theta:E\to M_{1}$ extending $\theta_{X}$. Define a linear order $\prec_{\zeta}$ on $E$ by $x\prec_{\zeta}y$ iff $\theta(x)\prec_{\beta}\theta(y)$. We have that $\prec_{\zeta}$ is a linear order on $E$ extending $\prec_{X}$ on $X$, and that $\theta:(E,\prec_{\zeta})\to(M_{1},\prec_{\beta})$ is a $\leq_{1}$-ordered graph embedding.
We may label the 11 copies of $C_{i}$ in $E$ as $C_{i,m}$ ($-5\leq m\leq 5$), with $\leq_{1}$-embeddings $\eta_{m}:C_{i}\to C_{i,m}\leq E$, and the corresponding copies of $d_{i+1}$ as $d_{i+1,m}\in C_{i,m}$, such that $d_{i+1,-5}\prec\cdots\prec d_{i+1,5}$ in $\prec_{\zeta}$. Let ${C_{i+1}}^{\prime}=C_{i,0}\cup\{d_{i+1,m}:-5\leq m\leq 5\}$. We have that $({C_{i+1}}^{\prime},\tau)\sqsubseteq_{s}(E,\tau)$, so ${C_{i+1}}^{\prime}\leq E$. So $\theta:({C_{i+1}}^{\prime},\prec_{\zeta})\to(M_{1},\prec_{\beta})$ is a $\leq_{1}$-ordered graph embedding.
We have that ${C_{i+1}}^{\prime}$ consists of a copy $C_{i,0}=\eta_{0}(C_{i})$ of $C_{i}$, where $\eta_{0}|_{X}=\operatorname{id}_{X}$ and $\eta_{0}|_{X}:(X,\prec_{X})\to(X,\prec_{\zeta})$ is order-preserving, together with witness vertices $d_{i+1,m}$ (where $1\leq|m|\leq 5$) for $d_{i+1,0}=\eta(d_{i+1})$.
Recall that $C_{i}$ consists of $D_{T}$ together with, for $1\leq j\leq i$, the witness vertices for $d_{j}$, and also that $X$ consists of $\operatorname{L}_{0}(D_{T})$ together with, for $1\leq j\leq i$, $d_{j}$ and its witness vertices.
Therefore $({C_{i+1}}^{\prime},\prec_{\zeta})$ consists of a graph-isomorphic copy $\eta_{0}(D_{T})$ of $D_{T}$, together with witness vertices in $\prec_{\zeta}$ for $\eta_{0}(d_{1})=d_{1},\cdots,\eta_{0}(d_{i})=d_{i}$ and witness vertices in $\prec_{\zeta}$ for an additional vertex $\eta_{0}(d_{i+1})$. We can therefore construct an ordered graph $C_{i+1}^{\prec}$ isomorphic to $({C_{i+1}}^{\prime},\prec_{\zeta})\in\mathcal{J}$ such that $C_{i+1}$ consists of $D_{T}$ together with witness vertices for $d_{j}$, $11\leq j\leq i+$ 1. This completes the induction step. We then let $C_{T}^{\prec}=C_{k}^{\prec}$. ∎
4.5. $(\mathcal{J},\leq_{1})$ does not have WAP
Proposition 4.8.
The class $(\mathcal{J},\leq_{1})$ does not have the weak amalgamation property.
Proof.
Suppose $(\mathcal{J},\leq_{1})$ has WAP, seeking a contradiction. Let $\{a_{0}\}\in\mathcal{J}$ be a singleton with the trivial linear order. Then there exists $\{a_{0}\}\leq_{1}A^{\prec}\in\mathcal{J}$ with $A^{\prec}$ witnessing WAP for $\{a_{0}\}$. Take $A\leq_{1}B\in\mathcal{C}_{1}$ witnessing for $A$ the expansion property of $\mathcal{J}$ over $(\mathcal{C}_{1},\leq_{1})$. (Here we use Theorem 2.30, recalling that $Y=X(\mathcal{J})$ is a minimal $G$-flow.)
Take $B^{+}\in\mathcal{D}_{1}$ such that the undirected reduct of $B^{+}$ is $B$. For each $v\in B^{+}$ with $\operatorname{d_{{}_{+}}\!}(v)=1$, add to $B^{+}$ a new vertex $v^{\prime}$ and out-edge $vv^{\prime}$, and call the resulting digraph $C\in\mathcal{D}_{1}$. Note that each vertex of $C$ has out-degree 00 or $2$. We have that $B\leq_{1}C^{-}$ as undirected graphs. Let $q$ be the maximum number of levels in any acyclic reorientation of $C$ (i.e. if $C$ when reoriented has levels $0,\cdots,n$, then $q=n+1$).
For $i=0,1$, let $C_{i}^{\prec}\in\mathcal{J}$ be $T_{i}(q)$-witness ordered graphs for $C$, using Lemma 4.7, and let $D_{i},Z_{i}$ denote $D_{T_{i}(q)},Z_{T_{i}(q)}$ (the notation here is introduced just above Definition 4.5).
As $B\leq C_{i}$ witnesses the expansion property for $A$, there exist $\leq_{1}$-ordered graph embeddings $\zeta_{i}:A^{\prec}\to(B,\prec_{C_{i}})\leq C_{i}^{\prec}$ ($i=0,1$). As $A^{\prec}$ witnesses WAP for $\{a_{0}\}$, there exists $D\leq M_{1}$ and $\leq_{1}$-ordered graph embeddings $\theta_{i}:C_{i}^{\prec}\to(D,\prec_{\beta})$ with $\theta_{0}\zeta_{0}(a)=\theta_{1}\zeta_{1}(a)$. By Lemma 4.6, $\theta_{i}|_{Z_{i}}:Z_{i}\to(D,\rho)$ are also digraph embeddings.
If $r$ is a vertex of $C$ of out-degree 00 in $C$, then as $\theta_{i}|_{Z_{i}}$ is a digraph embedding, $\theta_{i}(r)$ has out-degree 00 in $\theta_{i}(C)$. Also $\theta_{i}$ is a graph embedding, so preserves the sum of out-degrees, and as each vertex of $C$ has out-degree $2$ or 00 and $\theta_{i}(C)$ is $2$-oriented, we have that the vertices of $\theta_{i}(C)$ of out-degree $<2$ are exactly the $\theta_{i}(r)$ for $r$ a vertex of $C$ of out-degree 00.
Let $d=\theta_{i}\zeta_{i}(a)$, and let $U_{n}$ be the set of vertices of $(M_{1},\rho)$ that can be reached from $d$ by an outward-directed path of length $\leq n$. As the only vertices of $\theta_{i}(D_{i})$ of out-degree less than $2$ are the leaves of the copies of $T_{i}$, we have $U_{2q+1}\subseteq\theta_{i}(D_{i})$ ($i=0,1$).
We now obtain a contradiction by comparing the two cases $i=0$ and $i=1$. As $U_{2q+1}\subseteq\theta_{0}(D_{0})$, we have that $U_{2q+1}-U_{q-1}$ does not contain any (undirected) cycles. But as $U_{2q+1}\subseteq\theta_{1}(D_{1})$, we have that $U_{2q+1}-U_{q-1}$ contains a $4$-cycle - contradiction. ∎
This completes the proof of Theorem 4.1.
4.6. $\mathcal{LO}(M_{1})$ is not minimal
We now quickly show that $\mathcal{LO}(M_{1})$ is not in fact minimal itself.
Proposition 4.9.
$\mathcal{LO}(M_{1})$ is not a minimal flow.
Proof.
Let $\mathcal{Q}_{1}$ be the class of ordered graphs $A^{\prec}$ where $A\in\mathcal{C}_{1}$ and $\prec_{A}$ induces a $2$-orientation $\tau_{A}$ on $A$: that is $\tau_{A}=\{(x,y)\in E_{A}:y\prec_{A}x\}$ is a $2$-orientation (which must necessarily be acyclic, as $\prec_{A}$ is a linear order).
We will show that $\mathcal{Q}_{1}$ is a reasonable class of expansions of $(\mathcal{C}_{1},\leq_{1})$ (see Definition 2.21). Parts (2) and (3) of reasonableness are immediate. For parts (1) and (4), take $A^{\prec}\in\mathcal{Q}_{1}$ and $B\in\mathcal{C}_{1}$ with $A\leq_{1}B$ (where we allow $A^{\prec}=\varnothing$). Let $\tau_{A}$ be the acyclic $2$-orientation induced by $\prec_{A}$ on $A$. As $A\leq_{1}B$, there exists an acyclic $2$-orientation $\tau_{B}$ of $B$ extending $\tau_{A}$. Let $\prec_{0}=\{(b,b^{\prime})\in B^{2}:b\neq b^{\prime}$ and there exists an out-path from $b^{\prime}$ to $b$ in $\tau_{B}\}$. Then $\prec_{0}$ is a strict partial order on $B$. $\prec_{A}$ and $\prec_{0}$ are compatible, and so we may extend the partial order $\prec_{A}\cup\prec_{0}$ arbitrarily to a linear order $\prec_{B}$ on $B$. Then $\prec_{B}$ induces $\tau_{B}$, so $(B,\prec_{B})\in\mathcal{Q}_{1}$.
By Lemma 2.26, we therefore have that $X(\mathcal{Q}_{1})$ is a subflow of $\mathcal{LO}(M_{1})$. To see that it is a proper subflow, we produce a linear order on $M_{1}$ which does not induce an acyclic $2$-orientation. Let $\prec$ be the linear order of the Fraïssé limit of the order expansion $(\mathcal{C}_{1}^{\prec},\leq_{1})$ of $(\mathcal{C}_{1},\leq_{1})$. By genericity, there exists a graph $A\leq_{1}M_{1}$ consisting of vertices $a,b_{1},\cdots,b_{3}$ and edges $ab_{i}$ with $b_{i}\prec a$ ($1\leq i\leq 3$), so $\prec$ does not induce a 2-orientation. ∎
See [18, Section 4.6] for an explicit example of a minimal subflow of $\mathcal{LO}(M_{1})$.
Acknowledgements. The author would like to thank David Evans for his supervision during this project, which formed part of the second half of the author’s PhD thesis.